OpenStudy (anonymous):

rewrite the following expression as a single logarithm: 3log(x+4)+2log(x-7)-5log(x-2)-log(x^2)

OpenStudy (anonymous):

Remember the following 3 rules of logarithms: 1. Added logarithms can be combined to be a single logarithm of a product. 2. Subtracted logarithms can be combined to be a single logarithm of a fraction 3. Coefficients in front of logarithms can be taken as a power of the expression inside the logarithm.

OpenStudy (zzr0ck3r):

3log(x+4)+2log(x-7)-5log(x-2)-log(x^2) =log((x+4)^3)+log((x-7)^2)-log((x-2)^5)-log(x^2) = log(((x+4)^3(x-7)^2)/((x-2)^5*x^2)))

OpenStudy (anonymous):